Multiply 62 by 35, firstly by breaking down 35 in its terms (30 + 5) and secondly by breaking down 35 in its factors (5 x 7). Show all your steps.

(a) 62 x (30 + 5)

(b) 62 x (5 x 7)


1
Expert's answer
2022-05-24T18:10:00-0400

(a) 62*(30+5). No calculate it, we should remove the brackets. To remove the brackets in an algebraic expression, you multiply what is inside the brackets by what is outside them.

This is sometimes called expanding the brackets:

62*30+62*5=1860+310=2170

(b) 62*(5*7). In this problem we have product of two numbers in brackets multiplied by 62, so we can just remove brackets without multiplication 62 for each number in brackets:

62*5*7=310*7=2170